# OccamLLM: Fast and Exact Language Model Arithmetic in a Single Step

### v2

@@ 5 Discussion
We benchmarked our method on challenging arithmetic tasks, achieving 100% accuracy while GPT 4o achieves only 40% performance on average. We also benchmarked our method on mathematical problem solving tasks, demonstrating that the OccamLlama switch can accurately balance the LLM for reasoning and OccamNet for arithmetic, outperforming even GPT 4o and GPT 4o with Code Interpreter on tasks involving challenging arithmetic. Our work could enable smaller LLMs to be as performant as much larger LLMs in arithmetic. Moreover, integrating OccamLLM with larger LLMs like GPT 4o could further improve their arithmetic abilities without requiring a code interpreter, but we leave this for future work. Furthermore, at present, OccamLLM may not integrate with more advanced decoding techniques such as speculative decoding [38, 43]. We hope to explore these avenues in future work.
6 Broader Impact
We believe that, in addition to enabling fast, safe, and interpretable arithmetic, OccamLLM demonstrates a new paradigm for tool use. As a proof of concept for more complex tool use, we further train OccamLlama with a two layer Complete OccamNet with the primitives
@@ Appendix A Further Experiment Details and Results
For all training and evaluation runs, we used a single 32 GB NVIDIA Tesla V100 GPU. The LLM used for the OccamLLM system is Llama 3 8B Instruct.
In the experiments presented in Section 4, for each of the weight decoders and the switch, we used two-layer MLPs of input size 4096 (Llama 3 8B Instruct hidden size), intermediate size 64 and final size equal to the number of weights in the corresponding OccamNet layer or switch.
@@ A.1.1 OccamNet Decoder
To train the OccamNet decoder, we created a training dataset consisting of a 80,000 examples split in 40,000 single queries and 40,000 sequences of concatenated queries. In the first case, we sampled a single prompt of those described in 3.2 and formatted it using the Llama 3 8B Instruct chat template. In the second case, we concatenated multiple prompts described in 3.2 without the chat template.
@@
For the two-layer OccamNet run, we generated a large set of programmatically generated prompts of the form 3+974=3+97$\cdot$-4=3 + 97 ⋅ - 4 =, with the Llama 3 8B chat template applied.
@@ A.1.2 Switch Decoder
To train the switch decoder, we created a dataset of 50,000 examples. Each example consists of a series of variable length input-output pairs sampled from a collection of prompts, which are formatted in the Llama 3 8B Instruct chat template in a conversational style. For each pair, the tokens previous to the numbers that should be computed using OccamNet, which are the ones that the switch should not route to the LLM, are marked with a 1, and all the rest are marked with a 0.
25% of these pairs are created by taking one of the arithmetic expressions described above as input, and using as assistant output either “Answer = ”, the same expression repeated with an “=”, or nothing at all.
@@
70% of the pairs come from a collection of manually created and labeled examples, which illustrate in which cases the switch should route to OccamNet and, importantly, in which cases it shouldn’t. This collection was designed to cover a wide variety of situations where the LLM might need to use OccamNet for computations. Furthermore, it includes cases where the LLM should avoid calling OccamNet because doing so would produce a bad prediction. This is the case, for example, of instances where the LLM attempts to add three numbers simultaneously. If it were to use the 1-layer OccamNet, which can take 2 inputs at most, the result would be incorrect.
@@
The remaining 5% of the prompts come from multi-step reasoning problems. The output for these is not a full response, but only “The answer is ”. In such cases, a single-OccamNet-layer OccamLlama cannot solve the problem, so the switch output is set to 0.
@@ A.5 Experimental Results
Tables 4 and 5 show in more detail the accuracy of OccamLlama and other baselines on arithmetic and mathematical problem solving tasks. We measure accuracy as described in the main text.
We note here that on datasets with challenging arithmetic, such as Multiarith Float and MATH401, OccamLlama outperforms GPT 4o and GPT 4o Code. In fact, on MultiArith Float, OccamLlama 20 percentage points more accurate than GPT 4o and and 40 percentage points more accurate than Llama 3 8B. Similarly, on MATH401, OccamLlama is nearly 10 percentage points more accurate than GPT 4o and 25 percentage points more accurate than Llama 3 8B. Although MATH401 does not include word problems, it does include some arithmetic expressions that require multiple calls to OccamNet to solve, meaning it requires both reasoning (to determine how to break up the arithmetic expression) and arithmetic capabilities.
The only dataset on which OccamLlama struggles is GSM8K [40]. We believe the reason for this is primarily because this data involves a large number of fractions and percentages, which Llama does not convert to decimal format. As such, Llama often calls OccamNet with expressions such as “multiplying by 3/4 gives” or “5% of this gives.” Because the switch is not trained on many examples like these in which the number is not in decimal format, it does not realize that OccamNet cannot be used in these cases. Therefore, the switch triggers OccamNet, which is not capable of performing the correct operation because it requires two steps. Future work could address this issue by training the router with more data on this type of situation or by training an OccamLlama with a two layer OccamNet.
@@
Table 4: Accuracy on arithmetic tasks. Higher is better. Bold indicates best performance.
@@
Table 5: Accuracy on reasoning tasks. Higher is Better. Bold indicates best performance.
OccamLlama Llama 2 Llama 3 GPT 3.5 GPT 4o GPT 4o
7B Chat 8b Instruct Turbo Code
AddSub 91.4$\pm$1.4 78.0$\pm$2.1 93.4$\pm$1.2 95.4$\pm$1.1 97.0$\pm$0.9 97.5$\pm$1.1
GSM8K 68.8$\pm$1.3 36.0$\pm$1.3 79.8$\pm$1.1 84.8$\pm$1.0 96.1$\pm$0.5 94.0$\pm$1.7
MultiArith 99.5$\pm$0.3 76.0$\pm$1.7 99.8$\pm$0.2 97.2$\pm$0.7 99.7$\pm$0.2 99.5$\pm$0.5
MultiArith Float 97.5$\pm$0.6 23.3$\pm$1.7 57.3$\pm$2.0 77.3$\pm$1.7 96.2$\pm$0.8 89.5$\pm$2.2
MATH401 86.3$\pm$1.7 43.9$\pm$2.5 60.3$\pm$2.4 63.1$\pm$2.4 76.6$\pm$2.1 78.0$\pm$2.9
Single Eq 91.3$\pm$1.2 79.1$\pm$1.8 96.3$\pm$0.8 97.8$\pm$0.6 98.0$\pm$0.6 99.0$\pm$0.7
SVAMP 87.9$\pm$1.0 61.5$\pm$1.5 86.3$\pm$1.1 87.8$\pm$1.0 96.2$\pm$0.6 96.5$\pm$1.3
Average 85.8$\pm$0.5 53.3$\pm$0.7 82.1$\pm$0.6 86.5$\pm$0.5 95.2$\pm$0.3 93.4$\pm$0.7
Table 6: Relative error on arithmetic tasks (%). Lower is Better.
@@ B.2 MultiArith-Float
Appendix C Alternative Architectures and Losses
C.1 Alternative Architectures
@@ C.2 Alternative Losses
Appendix D Background on OccamNet
@@
    In Section D.1, we describe OccamNet’s architecture in more detail.
@@
    In Section D.2, we describe OccamNet’s sampling process.
@@
    In Section D.3, we describe OccamNet’s probability distribution.
@@
    In Section D.4, we describe OccamNet’s initialization process.
D.1 OccamNet Architecture
@@
In a standard OccamNet layer, each primitive is repeated exactly once in each layer. However, in Complete OccamNet, each primitive in the $l$th layer is repeated $A^{L-l}$ times, where $A$ is the maximum arity of the primitives. This is shown in Figure 6 in the transition from 6a to 6b. Complete OccamNet also concatenates each image layer to the next image layer, as shown in Figure 6c.
D.2 Sampling from OccamNet
@@
D.3 OccamNet’s Probability Distribution

### v4

@@ Appendix A Further Experiment Details and Results
For training and evaluation OccamLlama 8B, we used a single 32 GB NVIDIA Tesla V100 GPU. For OccamLlama 70B, we used two 80 GB NVIDIA A100 GPU.
In the experiments presented in Section 4, for each of the weight decoders and the switch, we used two-layer MLPs of input size 4096/8192 (Llama 3 8B/70B Instruct hidden size), intermediate size 64 and final size equal to the number of weights in the corresponding OccamNet layer or switch.
@@ A.1.1 OccamNet Decoder
To train the OccamNet decoder, we created a training dataset consisting of a 80,000 examples split in 40,000 single queries and 40,000 sequences of concatenated queries. In the first case, we sampled a single prompt of those described in 3.2 and formatted it using the Llama 3 Instruct chat template. In the second case, we concatenated multiple prompts described in 3.2 without the chat template.
@@
For the two-layer OccamNet run, we generated a large set of programmatically generated prompts of the form 3+974=3+97$\cdot$-4=3 + 97 ⋅ - 4 =, with the Llama 3 Instruct chat template applied.
@@ A.1.2 Switch Decoder
To train the switch decoder, we created a dataset of 50,000 examples (80,000 for OccamLlama 80B). For each example, the tokens previous to the numbers that should be computed using OccamNet, which are the ones that the switch should not route to the LLM, are labeled with a 1, and all the rest are labeled with a 0.
Half of the examples consist of a single prompt corresponding to a simple arithmetic expression as the ones described in Section 3.2. The token immediately at the beginning of the assistant response is labeled with a 1. Therefore, the trained system will answer directly to simple arithmetic queries that OccamNet can compute.
The remaining 25,000 examples consist each of a series of prompts which are formatted in the Llama 3 Instruct chat template in a conversational style. The input-output pairs used to create each sequence of prompts are distributed in the following way:
  • 25% of these pairs are created by taking one of the simple arithmetic expressions as input. The output is selected randomly between answering directly at the beginning of the assistant response, adding "Answer = " before the answer, or repeating the input expression before the answer. These examples train the switch to trigger OccamNet in different scenarios where the LLM needs to compute an answer.
@@
  • 70% of the pairs come from a collection of 43 manually created and labeled examples, which illustrate in which cases the switch should route to OccamNet and, importantly, in which cases it shouldn’t. This collection was designed to cover a wide variety of situations where the LLM might need to use OccamNet for computations. Furthermore, it includes cases where the LLM should avoid calling OccamNet because doing so would produce a bad prediction. This is the case, for example, of instances where the LLM attempts to add three numbers simultaneously. If it were to use the 1-layer OccamNet, which can take 2 inputs at most, the result would be incorrect.
@@
  • The remaining 5% of the prompts come from multi-step reasoning problems. We set the output for these not to a full response, but only “The answer is ”. In such cases, a single-layer OccamNet cannot compute the answer, so the output tokens are labeled with a 0. This trains the system to avoid routing to OccamNet when the later cannot compute the answer.
Below, we present some manually created examples and explain what case they cover. The tokens of the numbers in bold were marked with a 1 to signal that the switch should route to OccamNet, and the rest of the numbers and all of the text tokens were marked with a 0 to signal that OccamNet should not be called in these cases.
The cases we cover with these prompts are:
  1. 1.
    Single-layer OccamNet should be called to perform two-input arithmetic operations.
  2. 2.
    OccamNet should not be called to write numbers that already appeared and do not require further arithmetic.
  3. 3.
    OccamNet should not be called to round numbers.
  4. 4.
    OccamNet should not be called to perform inline operations nested within a larger expression because it might contradict the assumption that the last two numbers are involved in the operation.
  5. 5.
    OccamNet should not be used to convert magnitudes when the LLM does not explicitly write the conversion factor and relies on its internal knowledge; without the numeric value of the factor present in the text, OccamNet can’t give the correct result.
  6. 6.
    OccamNet can’t compute percentages, as this operation was not included in the training set of operations. It could be done in two steps.
  7. 7.
    Single-layer OccamNet can’t compute a fraction of a number, as this would require two operations.
  8. 8.
    Single-layer OccamNet can’t perform operations with more than 2 arguments.
User Query Assistant Response Cases
An author writes 7.6 pages per session. After 9 sessions, the total pages written are The author writes 7.6 pages per session. After 9 sessions, the author will have written 7.6 ×9 = 68.4 pages. The answer is 68. 1, 2, 3
I have 10 oranges and 6 apple trees, each of them with 3 apples. How much fruit do I have? The total number of pieces of fruit is 10 oranges + 6 trees times 3 apples = 10 + (6 ×3) = 10 + 18 = 28. The answer is 28. 1, 2, 4
Sally has 7 pennies and 8 nickels. How many pennies does she have in total? 1. Convert the number of nickels to pennies: 8 nickels is 40 pennies.
2. Add total number of pennies 40 + 7 = 47.
The answer is 47.
1, 2, 5
What is 84 percent of 230.77? 84% of 230.77 = 193.85
Explanation:
84 / 100 = 0.84
0.84 ×230.77 = 193.85
1, 2, 6
There were 39 cats on a rock. 3 boats came and carried away 5 cats each. How many cats were left? There were originally 39 cats. 3 boats came and each took away 5 cats. So 3 ×5 = 15.
39 - 15 = 24.
Then 3/7 of them ran away. 3/7 of 24 is 9.
24 - 9 = 15.
So there were 15 cats left.
1, 2, 7
11 people have 2 apples, 2 oranges, and 2 peaches each. How many pieces of fruit do they have? Each person has 2 + 2 + 2 = 6 pieces of fruit.
In total, they have 6 ×11 = 66.
The answer is 66.
1, 2, 8
@@ A.5 Experimental Results
Tables 5 and 6 show in more detail the accuracy of OccamLlama and other baselines on arithmetic and mathematical problem solving tasks. We measure accuracy as described in the main text.
We note here that on datasets with challenging arithmetic, in particular Multiarith Float and MATH401, OccamLlama 8B outperforms even GPT 4o and GPT 4o Code. In fact, on MultiArith Float, OccamLlama 8B is nearly 10 percentage points more accurate than GPT 4o + Code and and more than 40 percentage points more accurate than Llama 3 8B. Similarly, on MATH401, OccamLlama 8B is 7 percentage points more accurate than GPT 4o + Code and nearly 25 percentage points more accurate than Llama 3 8B. Although MATH401 does not include word problems, it does include some arithmetic expressions that require multiple calls to OccamNet to solve, meaning it requires both reasoning (to determine how to break up the arithmetic expression) and arithmetic capabilities.
The only datasets on which OccamLlama 8B performs substantially worse than Llama 3 8B are GSM8K [40] and Single Eq [41]. We believe a contributor to this is that these datasets include many problems that involve either fractions and percentages, which Llama does not convert to decimal format, or equations with unknown variables. As such, Llama often calls OccamNet with expressions such as “multiplying by 3/4 gives,” “5% of this gives,” or “adding 5 to both sides of x-5 = 11 gives.” Because the switch is not trained on many examples like these in which the number is not in decimal format, it does not realize that OccamNet should not be used in these cases. Therefore, the switch triggers OccamNet, which is not capable of performing the correct operation (these types of operations are not acheivable with a 1-layer OccamNet). Future work could address this issue by training the switch with more data on this type of situation or by training an OccamLlama with a two layer OccamNet.
Finally, as noted in the main text, OccamLlama 70B achieves significant performance improvement over OccamLlama 8B across a number of benchmarks and outperforms GPT 4o and GPT 4o + Code on average. This demonstrates that OccamLLM improves with the base language model and suggests that combining OccamLLM with more capable models such as GPT 4o could be a promising avenue for future research.
@@
Table 5: Percent accuracy on arithmetic tasks. Higher is Better. Bold indicates best performance.
@@
Table 6: Percent accuracy on reasoning tasks. Higher is Better. Bold indicates best performance.
OccamLlama OccamLlama Llama 3 Llama 3 GPT 3.5 GPT 4o GPT 4o
8B 70B 8b Instruct 70B Instruct Turbo Code
AddSub 91.6$\pm$1.4 96.5$\pm$0.9 93.4$\pm$1.2 97.2$\pm$0.8 95.4$\pm$1.1 97.0$\pm$0.9 97.5$\pm$1.1
GSM8K 73.5$\pm$1.2 90.1$\pm$0.8 79.8$\pm$1.1 94.8$\pm$0.6 84.8$\pm$1.0 96.1$\pm$0.5 94.0$\pm$1.7
MultiArith 99.2$\pm$0.4 98.5$\pm$0.5 99.8$\pm$0.2 98.2$\pm$0.5 97.2$\pm$0.7 99.7$\pm$0.2 99.5$\pm$0.5
MultiArith Float 98.2$\pm$0.5 97.7$\pm$0.6 57.3$\pm$2.0 76.3$\pm$1.7 77.3$\pm$1.7 96.2$\pm$0.8 89.5$\pm$2.2
MATH401 85.0$\pm$1.8 89.5$\pm$1.5 60.3$\pm$2.4 71.6$\pm$2.3 63.1$\pm$2.4 76.6$\pm$2.1 78.0$\pm$2.9
Single Eq 92.9$\pm$1.1 96.9$\pm$0.8 96.3$\pm$0.8 97.6$\pm$0.7 97.8$\pm$0.6 98.0$\pm$0.6 99.0$\pm$0.7
SVAMP 88.6$\pm$1.0 93.2$\pm$0.8 86.3$\pm$1.1 94.5$\pm$0.7 87.8$\pm$1.0 96.2$\pm$0.6 96.5$\pm$1.3
Average 89.9$\pm$1.2 94.6$\pm$0.9 81.9$\pm$1.5 90.0$\pm$1.2 86.2$\pm$1.4 94.2$\pm$1.0 93.4$\pm$1.7
Table 7: Relative error (%) on arithmetic tasks. Lower is Better. Bold indicates best performance.
@@ Appendix B Example OccamLlama Generations
We also include three miscellaneous prompts and OccamLlama 8B generations to demonstrate OccamLlama’s ability to handle more challenging prompts.
@@ B.2 MultiArith-Float
B.3 Miscellaneous prompts
These miscellaneous prompts demonstrate OccamLlama’s ability to handle more challenging prompts. Examples 1 and 2 show that OccamLlama can handle prompts where the numbers are provided as words and not as decimals. This works because we can choose to parse strings such as “three” and “four” as 3 and 4, respectively, when parsing inputs to OccamNet. Example 3 shows that OccamLlama can handle prompts in which many irrelevant numbers are included. This works because OccamLlama uses Llama to reason and select the correct numbers to use before feeding these numbers into OccamNet.
Example 1
Prompt:
Six minus seven =?
Response:
-1. Six minus seven is equal to -1.
Example 2
Prompt:
Three + four = ?
Response:
7. 3 + 4 = 7. Let me know if you have any other questions!
Example 3
Prompt:
In 1932, 100 planets were discovered and in 1933, another 200 were found. What is the total number of planets discovered between 1932 and 1933?
Response:
The number of planets discovered in 1932 was 100. The number of planets discovered in 1933 was 200. So the total number of planets discovered is 100 + 200 = 300. 300.
Appendix C OccamLLM Robustness Experiments
We find that OccamLlama displays remarkable generalization capabilities on out-of-distribution problems. To demonstrate this, we show below two out-of-distribution tasks on which OccamLlama performs remarkably well.
C.1 Non-textual Training
First, we train the OccamNet decoder from scratch, using only numeric expressions and absolutely no text at all. This means that any problem with text, such as a word problem, is far out-of-distribution of the OccamNet decoder’s training data. We test this model (using the standard router), which we denote OccamLlama 8B Arith, on the mathematical reasoning benchmarks and obtain remarkably good results, shown in Table 8.
Table 8: Accuracy on reasoning tasks. Higher is Better.
OccamLlama OccamLlama Llama 2 Llama 3 GPT 3.5 GPT 4o GPT 4o
8B 8B Arith 7B Chat 8b Instruct Turbo Code
AddSub 91.6$\pm$1.4 92.7$\pm$1.3 78.0$\pm$2.1 93.4$\pm$1.2 95.4$\pm$1.1 97.0$\pm$0.9 97.5$\pm$1.1
GSM8K 73.5$\pm$1.2 71.6$\pm$1.2 36.0$\pm$1.3 79.8$\pm$1.1 84.8$\pm$1.0 96.1$\pm$0.5 94.0$\pm$1.7
MultiArith 99.2$\pm$0.4 98.5$\pm$0.5 76.0$\pm$1.7 99.8$\pm$0.2 97.2$\pm$0.7 99.7$\pm$0.2 99.5$\pm$0.5
MultiArith Float 98.2$\pm$0.5 95.3$\pm$0.9 23.3$\pm$1.7 57.3$\pm$2.0 77.3$\pm$1.7 96.2$\pm$0.8 89.5$\pm$2.2
MATH401 85.0$\pm$1.8 85.8$\pm$1.7 43.9$\pm$2.5 60.3$\pm$2.4 63.1$\pm$2.4 76.6$\pm$2.1 78.0$\pm$2.9
Single Eq 92.9$\pm$1.1 92.1$\pm$1.2 79.1$\pm$1.8 96.3$\pm$0.8 97.8$\pm$0.6 98.0$\pm$0.6 99.0$\pm$0.7
SVAMP 88.6$\pm$1.0 88.8$\pm$1.0 61.5$\pm$1.5 86.3$\pm$1.1 87.8$\pm$1.0 96.2$\pm$0.6 96.5$\pm$1.3
Average 89.9$\pm$1.1 89.3$\pm$1.1 56.8$\pm$1.8 81.9$\pm$1.3 86.2$\pm$1.2 94.2$\pm$0.8 93.4$\pm$1.5
The OccamLlama 8B Arith performs on par with the model trained with both numbers and text, even achieving higher accuracy on some benchmarks. This shows that the OccamLLM framework is robust, and points towards the fact that the representations of arithmetic that are built in the transformer body of the LLM and extracted by the OccamLLM Decoder are very general.
In contrast, we expect that finetuning Llama to perform arithmetic using only numeric examples and no text whatsoever would lead to extreme catastrophic forgetting and poor arithmetic performance on word problems. As such, we believe this data shows a remarkable generalization and robustness of OccamLLM.
C.2 Multilingual Reasoning
To further demonstrate OccamLlama’s generalization capabilities and also show that OccamLlama can handle non-English generation, we tested OccamLlama on the Multilingual Grade School Math Benchmark (MGSM) [44], a dataset consisting of GSM8K translated into 10 languages (Bengali, Chinese, French, German, Japanese, Russian, Spanish, Swahili, Telugu, and Thai). For these experiments, we prompted the LLMs to write their answers in the same language as the problem statement. Otherwise, the LLM would typically respond always in English, defeating the purpose of the experiment. We compute the drop in accuracy when switching from English to another language, given by the accuracy of a model on the English dataset minus the accuracy of the model on the dataset in a given language. The results are shown in Figure 5.
The table above shows that OccamLlama and Llama have similar performance drops between the English dataset and each non-English language dataset. On most languages and on average, OccamLlama has a smaller performance drop than Llama. The fact that OccamLlama (the decoders for which have never been trained on other languages) has a smaller average out-of-distribution performance drop than Llama (a model trained on over 750 billion tokens of non-English text) is in our opinion quite remarkable.
We believe that this test demonstrates OccamLlama’s ability to handle many languages and its robustness against out-of-distribution data.
Figure 5: Model performance degradation for each language relative to English in the MGSM dataset. OccamLlama 8B’s performance degradation is considerably less than Llama 3 8B’s performance degradation, demonstrating strong multilingual and generalization capabilities.
Appendix D Alternative Architectures and Losses
D.1 Alternative Architectures
@@ D.2 Alternative Losses
Appendix E Background on OccamNet
@@
    In Section E.1, we describe OccamNet’s architecture in more detail.
@@
    In Section E.2, we describe OccamNet’s sampling process.
@@
    In Section E.3, we describe OccamNet’s probability distribution.
@@
    In Section E.4, we describe OccamNet’s initialization process.
E.1 OccamNet Architecture
@@
In a standard OccamNet layer, each primitive is repeated exactly once in each layer. However, in Complete OccamNet, each primitive in the $l$th layer is repeated $A^{L-l}$ times, where $A$ is the maximum arity of the primitives. This is shown in Figure 7 in the transition from 7a to 7b. Complete OccamNet also concatenates each image layer to the next image layer, as shown in Figure 7c.
E.2 Sampling from OccamNet
@@
E.3 OccamNet’s Probability Distribution